OpenStudy (anonymous):

Find S25 for for the following arithmetic sequence -7+ -2 + 3 + 8 +… I did the problem and got 113. Is that right?

OpenStudy (raden):

for to find the value a25 = a1+(25-1)5 = -7+120=113, u are right. but i think u want to find the value S25 not a25 right ?

OpenStudy (raden):

therefore, use the formula : Sn = n/2 * (a1 + an) S25 = 25/2 * (-7 + 113) = ...

OpenStudy (anonymous):

Yeah I need to find the value of S25

OpenStudy (raden):

well, recalculate it ^